Suppose the cost function for a firm is given byC(Q) = 60 + 4Q+ 2Q2.(a)If the firm sells output in a perfectly competitive market and other firms in the industry sell output at a price of $20, what level of output should the firm produce to maximize profits or minimize losses?(b)What will be the level of profits or losses if the firm makes the optimal decision?(c)After you’ve worked the problem, reconsider the original situation (with price of $20) but with fixed costs equal to 20. Finally reconsider the original situation (with fixed costs equal to $60) but with a market price of $8.
Answer #1

ANSCOTR Cost function CCQ) = 60 + 40-1202 (a) Price: $200 Marginal cost: 4+ 40. price is So equal profit to mopo mizing condi(c) when foxed cost- 20. Total Cost CC 200 401 202 4+ 40. Marginal cost Price $200 P= MC 4t4Q=20 40 = 16 0 to 4 умдi rnfuЯ оwhen fired Costa $60 Tota cost cco) - bor 40+ 202 Marginal Costa 4490 pace = $8. р мс whQzde 407 4 = unit us profot maki mizi